OpenStudy (anonymous):

Five more than a number tripled is half the number. What is the number? A. 6 B. –2 C. –1 D. 1

OpenStudy (johnweldon1993):

Write out the statements... 5 more than = + 5 A number tripled = 3n is half the number = n/2 Put it all together... 5 more than a number tripled is half the number = 3n + 5 = n/2 Now simplify and solve for 'n'

OpenStudy (anonymous):

-2 so b